OpenStudy (anonymous):

find the sum or difference. simplify if possible. 1/3 + 2/9 =

OpenStudy (anonymous):

help

Directrix (directrix):

1/3 + 2/9 = 9 is a common denominator for 3 and 9 1/3 = 3/9 and 2/9 = 2/9 So, 1/3 + 2/9 = 3/9 + 2/9 = 5/9.